Status Report

ISS On-Orbit Status 8 May 2002

By SpaceRef Editor
May 8, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted previously
or below. Day 552 of continuous station occupancy.

On wake-up (2:00 am EDT), the crew found kudos from MCC-H for their
excellent performance on yesterday’s “Good Morning America” TV event, while
POC (Payload Operations Center) particularly rejoiced over Dan Bursch’s
observing new wheat seedlings in BPS plant growth chamber #1.

Before breakfast, all ISS crewmembers completed the periodic Russian medical
experiment protocols of calf volume measurement (PZEh-MO-7) and body mass
measurement (PZEh-MO-8). FE-2 Bursch then closed out the experiments and
stowed the equipment. [Calf measurements (left leg only) are taken with the
ISOG device, a custom-sewn fabric cuff that fits over the calf, using the
knee and lower foot as fixed reference points, to provide a rough index of
deconditioning in zero-G and effectiveness of countermeasures. For
determining body mass in zero-G (where things are “weightless” but not
massless), Russian IM “scales” measure the inertial forces that arise during
the oscillatory motion of a mass driven by two helical metering springs with
known spring constants. By measuring the time period of each oscillation of
the unknown mass (the crewmember) and comparing it to the period of a known
mass, the crewmember’s mass is calculated by the computer and displayed]

CDR Yuri Onufrienko performed an additional Russian MedOps assessment on the
TVIS treadmill, called PROFILAKTIKA. [The test is analogous to the MO-3
fitness test, which calls for a gradually accelerating walk-jog-run up to 8
km/h speed within five minutes, followed by a gradual slow-down. After the
run, the heart rate at rest is also recorded for three more minutes.
PROFILAKTIKA adds to this a gas analysis, blood tests for lactate and
Creatine Kinase and a subjective assessment by Yuri of activity difficulty.
Today’s test focused on the ECG (electrocardiogram) part of the assessment,
while the blood analysis will be done in part 2 tomorrow.]

The crew completed an emergency evacuation drill/OBT (on-board training)
designed to refresh procedures and escape route in the event of cabin
depressurization. [Objectives of the periodic exercise are to review depress
response procedures and all incorporated actions to locate and isolate the
leaking compartment, practice crew coordination necessary to perform
emergency depress response measures, and demonstrate familiarity with
support equipment such as vacuum gauge and depress nomograms, hatch
operations, etc. used in depress response procedures. Starting from a cabin
pressure warning at 720 mmHg, the crew has about 45 minutes for a number of
major actions (searching for leak, closing hatches, removing equipment,
monitoring pressure) during which the pressure drops to 695 mmHg (simulated
on laptop display).]

Subsequently, CDR Onufrienko worked on the rotational hand controller
(RUO-2) of the TORU teleoperated rendezvous/docking system in the SM,
tightening its mounting which had exhibited excessive free play during
recent tests.

At about 6:00 am EDT, FE-1 Carl Walz terminated the successful regeneration
of the EMU Metox canisters in the Airlock (A/L) bakeout oven, started
yesterday (15.5 hours earlier). Later in the day, he prepared the
regeneration of the second batch of Metox canisters, to start at 4:30 pm. As
suggested by Moscow, the crew was asked to check on the system at regen
start +10 hours, i.e., at 2:30 am tomorrow morning (time of greatest heat),
and to report on what they see or smell. At that time, they are also to take
a reading with the CSA-CP (compound specific analyzer-combustion products).

Onufrienko’s major activity today, with 2.5 hrs reserved on the timeline,
was preventive maintenance on the SM ventilation subsystem (SV). This
consists of periodic filter replacements for both groups (A, B), followed by
a thorough cleaning of selected screens and grills on the SKV-1/SKV-2
gas-liquid heat exchanger, gas analyzer (GA) ventilation, fan screens, and
the flexible air duct. Later, he also replaced filters in the FGB’s dust
collectors (PS1, PS2) and cleaned the VD1 and VD2 air ducts in the DC-12
docking module with the vacuum cleaner.

Carl Walz meanwhile completed troubleshooting and recovery of the failed
MCOR (medium-rate communications outage recorder), i.e., powering it on and
working on its laptop to check on proper boot-up and then investigate the
cause of the shutdown of the important unit. Main thrust of the
troubleshooting was to change the MCOR temperature configuration file by
raising a temp limit setting from 40 deg C to 47 deg, the suspected cause of
the shutdown. By late afternoon, MCOR was reported to be operating.

FE-2 Dan Bursch worked with the TOCA (total organic carbon analyzer),
deploying it in the SM and mixing its reagent solution, then flushing the
system with deionized water via a syringe. Sample analysis for organic
carbon will be performed tomorrow (5/9).

Bursch continued BPS (biomass production system) activities, today
collecting water samples from the NDS (nutrient delivery system) and the HCS
(humidity control system) reservoirs. [While the BPS PIs are very happy with
the data and plant growth to date, the humidity control in PGC2 (plant
growth chamber 2) has been turned off and water has accumulated. The BPS
team does not believe that PGC2 HCS can be re-enabled without endangering
the other three chambers, and the plan is to have Bursch move the wheat root
module in PGC2 to PGC4 on 5/10, swapping places with the Brassica plants.
The wheat plant experiment requires a functioning HCS, whereas the Brassica
experiment is not as dependent on humidity control.]

Dan also was scheduled to do a reboot of the ER4 RIC (EXPRESS Rack 4/rack
interface controller), but since a recent as-yet-unexplained failure of the
PL1 MDM (payload computer #1) occurred during a ground-commanded ER4 reboot
attempt, some steps were to be taken beforehand to capture PL MDM data in
the event of a failure and to prevent the C&C MDM from deactivating the PL-2
MDM if it fails to mode to Diagnostics as it did the previous time. By doing
this, MCC-H hopes to take more data dumps as well as preserve the PL MDM
event log, which could be useful in diagnosing the cause of the failure.

Bursch completed another EVARM (EVA radiation monitoring) badge reading,
which meets the EVA-minus-28-days requirement.

Walz completed his weekly session with the NTXN interpersonal “Interactions”
experiment, recording mood, group and journal observations in a
privacy-protected (encrypted) database (not downlinked but eventually
carried down on its hard disk drive).

Daily routine tasks of SOSH life support systems maintenance and Lab payload
status checks were completed by Carl, while Dan prepared the daily “delta”
file to keep the IMS up to date. Walz also continued preparing return items
for the upcoming UF-2/STS-111 flight with the Expedition 5 crew (still
scheduled for launch on 5/30, docking on 6/1 and landing on 6/11).

The Elektron oxygen generator is still failed, now with both subsets, and
work continues to recover the system.

Today’s target areas for CEO (crew earth observation) were Cyclone —
Arabian Sea (Dynamic Event: Tropical cyclone 01A continues to move toward
the Arabian coast — the first major storm of the season. Using the ESC
[electronic still camera], the crew was to document internal structures
[bands, etc] of this increasingly compact storm. Their photos would fulfill
a request from the NASA Earth Observatory website for current images of this
storm), L. Nasser, Toshka Lakes; Egypt (Toshka Lakes were on track; sun
angle was high, so slight underexposure of this bright landscape should
reveal details of new roads, other construction projects around the new
lakes), Rift Triple Junction, Ethiopia (the black shield-shaped volcanic
complex of Erta Ale passed directly beneath the ISS. Some of the oldest
Early Man fossils have been found just east of the volcanic vents; detailed
photos of this area will contribute to archaeological investigations in the
area. Of interest: a stereo photo mapping strip of the network of NW-, NE-
and SW-oriented faults that offset the dark lava flows of the depression),
Somalia Coast (crew to look both left and right of track as you cross the
Somalian coastal plain and document the extent of dune fields and condition
of the vegetation), Tropical Cyclone Kesiny (Dynamic Event: Of interest:
using the ESC to document the movement of tropical cyclone Kesiny toward
northern Madagascar. The storm center was right of track and is expected to
reach hurricane force by Thursday [5/9]. ISS photos would fulfill a request
from the NASA Earth Observatory website for current images of this storm),
Rukwa Transform, Tanzania (weather conditions appeared uncommonly favorable
for taking a detailed stereo photo mapping strip from the central Lake
Tanganyika rift valley, through its junction with the Lake Rukwa transform
fault valley, then past the Rungwe volcanoes and along the upper Lake Malawi
rift valley. Crew to focus on fault escarpments and on the volcanic vents at
the north end of L. Malawi), Central American-Cuban Smoke (Dynamic Event:
ESC images have also been requested of the smoke and fires in Central
America and Cuba. Downlinked digital images of this situation will be
forwarded for publication on the NASA Earth Observatory website. ISS track
essentially bisected the Gulf of Mexico, from New Orleans to the west tip of
Cuba. To left of track, crew was to document smoke masses moving from the
Yucatan Peninsula, Honduras and Nicaragua, northward over the western Gulf
of Mexico. To right of track, they were to track the movement of smoke from
Cuban fires).

U.S. and Russian Segment Status (as of 2:00 pm EDT):

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is On, operating on 32 Amps. Vozdukh CO2 scrubber is
    ON in MANUAL cycle mode #5 (vacuum pump failed). U.S. CDRA CO2 scrubber is
    Off. BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode, bed #2 in
    Purify mode.

  • SM Working Compartment: Pressure (mmHg) — 754, temperature (deg C) — 26.1,
    ppO2 (mmHg) — 150.9, ppCO2 (mmHg) — 2.5.

  • SM Transfer Compartment: Pressure (mmHg) — 762, temperature (deg C) —

  • FGB Cabin: Pressure (mmHg) — 756, temperature (deg C) — 19.3.

  • Node: Pressure (mmHg) — 747.71, temperature (deg C) — 21.6 (shell); ppO2
    (mmHg) — n/a; ppCO2 (mmHg) — n/a.

  • U.S. Lab: Pressure (mmHg) — 749.66, temperature (deg C) — 24.1, ppO2
    (mmHg) — n/a; ppCO2 (mmHg) — n/a;

  • Joint Airlock (Equip. Lock): Pressure (mmHg) — 749.46, temperature (deg
    C) — 21.4;hell heater temp (deg C) — 22.7, ppO2 (mmHg) — n/a; ppCO2
    (mmHg) — n/a.

  • PMA-1: Shell heater temp (deg C) — 22.2

  • PMA-2: Shell heater temp (deg C) — 18.9

(Note: Partial pressures ppO2 and ppCO2 in U.S. segment [USOS] not available
because MCA [major constituent analyzer] is failed and in Extended Life mode
[= a state that preserves mass spectrometer vacuum but produces no pp
data]). MSA (mass spectrometer assembly) and VGA (verification gas assembly)
have been removed for return to Earth.

Electrical Power Systems (EPS):

  • Beta Gimbal Assembly (BGA) 2B in Autotrack mode, BGA 4B in Autotrack mode (dual-angle test ended at 9:00 am EDT).
  • SM batteries: Battery #6 is offline (failed); battery #5 is in cycle mode;
    all other batteries (6) are in “partial charge” mode.
  • FGB: Battery #1 is offline (storage mode); all other batteries (5) are in
    “partial charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Thermal Control Systems:

  • Air conditioner SKV-1 is On. SKV-2 is Off.

Command & Data Handling Systems:

  • C&C-3 MDM is prime, C&C-2 is back-up, and C&C-1 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is operational; PL-2 MDM on Standby.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational.

Communications Systems:

  • All Russian communications & tracking systems are nominal.
  • S-band is operating nominally.
  • Ku-band is operating nominally.
  • Audio subsystem operating nominally.
  • Video subsystem operating nominally.
  • MCOR (medium-rate communications outage recorder) is operational. Recovery
  • attempt today appears to have been successful.


  • SSRMS/Canadarm2 at Progress viewing position, with Keep Alive power on both strings.
  • RWS (robotics workstations) are Off.
  • Troubleshooting continues on SSRMS Prime string Wrist Roll (WR) joint.

ISS Orbit (as of this morning, 9:03 am EDT):

  • Mean altitude — 390.0 km
  • Apogee — 395.4 km
  • Perigee — 384.6 km
  • Period — 92.4 min.
  • Inclination (to Equator) — 51.64 deg
  • Eccentricity — 0.0007936
  • Orbits per 24-hr. day — 15.59
  • Altitude decrease 350 m (mean) in last 24 hours
  • Solar Beta Angle: -9 deg (magnitude decreasing)
  • Revolutions since FGB/Zarya launch (Nov. ’98) — 19784
  • Current Flight Attitude — LVLH (local vertical/local horizontal =
  • “earth-fixed”: z-axis in local vertical, x-axis in velocity vector [yaw: -10
  • deg, pitch: -7.25 deg., roll: 0 deg]).

For more on ISS orbit and naked-eye visibility dates/times, see

SpaceRef staff editor.